You have recently finished your MBA at the DarnIt Business School. Naturally, you must purchase a new BMW immediately. The car costs about $21,000. The bank quotes an interest rate of 15% APR for a 72 month loan with a 10 percent down payment. You plan on trading the car in for a new one in two years.

1)What is the effective annual interest rate be on the loan?
Answer #1

EAR = (1+r)^n - 1

r is int rate per month and n is no. of months per anum

= ( 1 + 0.00833)^12 - 1

= 1.00833^12 - 1

= 1.1047 - 1

= 0.1047 i.e 10.47%